OpenStudy (anonymous):

what # is 120% is 70

hero (hero):

is/of = %/100 x/70 = 120/100 100x = 8400 x = 8400/100 x = 84

hero (hero):

By the way, you wrote the question wrong. It should say, "What number is 120% of 70".

OpenStudy (anonymous):

# or number its the same thing

hero (hero):

That's not what I'm talking about

OpenStudy (anonymous):

oh yeah sorry my bad